Cute Camel Woodcraft Company is a small firm, and several of its managers are worried about how soon the firm will be able to recover its initial investment from Project Sigmas expected future cash flows. To answer this question, Cute Camels CFO has asked that you compute the projects payback period using the following expected net cash flows and assuming that the cash flows are received evenly throughout each year Complete the following table and compute the projects conventional payback period. For full credit, complete the entire table Year O Year 1 Year 2 Year 3 Expected cash flow Cumulative cash flow -5,000,000 $2,000,000 $4,250,000 $1,750,000 Conventional payback period The conventional payback period ignores the time value of money, and this concerns Cute Camels CFO. He has now asked you to compute Sigmas discounted payback period, assuming the company has a 10% cost of capital Complete the following table and perform any necessary calculations. Round the discounted cash flow values to the nearest whole dollar, and the discounted payback period to the nearest two decimal places. For full credit, complete the entire table Year O Year 1 Year 2 Year 3 Cash floww Discounted cash flow Cumulative discounted cash flow -5,000,000 $2,000,000 $4,250,000 $1,750,000 Discounted payback periodWhich version of a projects payback period should the CFO use when evaluating Project Sigma, given its theoretical superiority? O The regular payback period O The discounted payback period One theoretical disadvantage of both payback methods-compared to the net present value method-is that they fail to consider the value of the cash flows beyond the point in time equal to the payback period. How much value does the discounted payback per od method fa to recognize due to this theoretical deficency? O $1,645,380 O $3,132,983 $4,827,198 O $1,314,801
Answer #1

Ans: 1.

Year Cash flows Cummulative cash flows
0 -5000000 -5000000
1 2000000 -3000000
2 4250000 1250000
3 1750000 3000000

Conventional payback period = 1 + [3000000 / 4250000]

= 1 + 0.70588

= 1.70588

2.

Year Cash flows (A) Discount rate( Note 1) (B) Discounted cash flow a(A*B) Cummulative Discounted cash flow
0 -5000000 1 -5000000 -5000000
1 2000000 0.909090909 1818181.818 -3181818.182
2 4250000 0.826446281 3512396.694 330578.5124
3 1750000 0.751314801 1314800.902 1645379.414

Discounted Payback period = 1 + ( 3181818.182 / 3512396.694 )

= 1 + 0.90588

= 1.90588

3. CFO considers Discounted payback period as a superiority

4. Amount of value discounted payback period did not recognize = $ 1645379.414

Cummulative balance in discounted payback period.

Note 1 :

r = rate of interest , n = number of years

Present value (1+r)^n Discount rate( Note 1) (B)
1/(1+10%) 0.909090909
1/(1+10%)^2 0.826446281
1/(1+10%)^3 0.751314801